OpenStudy (anonymous):

The total profit (in dollars) from the sale of x espresso machines is P(x)=30x−0.7x^2−290 Evaluate the marginal profit at the following values: (A) P'(200) = (B) P'(350) =

OpenStudy (anonymous):

differentiate P(x) and you re good to go

OpenStudy (anonymous):

\[P'(x)=30-0.7(2)x\]

OpenStudy (anonymous):

\[P'(x)=30-1.4x\]

OpenStudy (anonymous):

from there do i just plug in the value of x right

OpenStudy (anonymous):

\[P'(200)=30-1.4(200)\]

OpenStudy (anonymous):

yeah same like this

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

yw